Equipment was acquired at the beginning of the year at a cost of $75,000. The equipment was depreciated using the straight-line method based on an estimated useful life of six years and an estimated residual value of $7,740.

Required:

(a) What was the depreciation expense for the first year?
(b) Assuming the equipment was sold at the end of the second year for $56,786, determine the gain or loss on sale of the equipment.
(c) Journalize the entry to record the sale. Refer to the Chart of Accounts for exact wording of account titles.
Answer #1

(a) Depreciation expense for the first year = (Cost - Salvage value) / Estimated useful life

= ($75,000 - 7,740) / 6 years

= $11,210

(b) Gain on sale of the equipment at the end of the second year Calculation:

Book value at the end of second year = Original cost - Two years depreciation

= 75,000 - (11,210+11,210)

= 75,000 - 22,420

= $52,580

Therefore, Gain on sale of the equipment at the end of the second year = Sale value - Book value at the end of second year

= 56,786 - 52,580

= $4,206

(C)

Journal entries
GENERAL JOURNAL DEBIT ($) CREDIT ($)
Cash          56,786.00
Accumulated Depreciation of Equipment (11,210 +11,210)          22,420.00
Equipment          75,000.00
Gain on Disposal of Equipment            4,206.00
(To record the sale of Equipment)